### v2

# Mixture-of-Skills: Learning to Optimize Data Usage for Fine-Tuning Large Language Models

Abstract
Large language models (LLMs) are typically fine-tuned on diverse and extensive datasets sourced from various origins to develop a comprehensive range of skills, such as writing, reasoning, chatting, coding, and more. Each skill has unique characteristics, and these datasets are often heterogeneous and imbalanced, making the fine-tuning process highly challenging. Balancing the development of each skill while ensuring the model maintains its overall performance requires sophisticated techniques and careful dataset curation. In this work, we propose a general, model-agnostic, reinforcement learning framework, Mixture-of-Skills (MoS), that learns to optimize data usage automatically during the fine-tuning process. This framework ensures the optimal comprehensive skill development of LLMs by dynamically adjusting the focus on different datasets based on their current learning state. To validate the effectiveness of MoS, we conduct extensive experiments using three diverse LLM backbones on two widely used benchmarks and demonstrate that MoS substantially enhances model performance. Building on the success of MoS, we propose MoSpec, an adaptation for task-specific fine-tuning, which harnesses the utilities of various datasets for a specific purpose. Our work underlines the significance of dataset rebalancing and present MoS as a powerful, general solution for optimizing data usage in the fine-tuning of LLMs for various purposes.
Mixture-of-Skills: Learning to Optimize Data Usage for Fine-Tuning Large Language Models
[…]
1 Introduction
Large language models (LLMs) have demonstrated their extraordinary capabilities and are expected to proficiently master a diverse range of skills (Ouyang et al., 2022; Sanh et al., 2022; OpenAI, 2023; Anil et al., 2023b; Touvron et al., 2023a, b; Anil et al., 2023a; Mesnard et al., 2024), such as writing, reasoning, chatting, coding, and more, through supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) on an extensive collection of datasets from various sources (Bai et al., 2022; Longpre et al., 2023; Ding et al., 2023; Wang et al., 2024b). Each dataset contributes unique elements to the model’s skill set, but this diversity also brings challenges.
Training Data $D_{\textrm{trn}}$corer $\boldsymbol{\psi}$LLM $\boldsymbol{\theta}$$\tilde{i}\sim p_{\boldsymbol{\psi}}(N)$$(\boldsymbol{x},\boldsymbol{y})\sim D_{\textrm{trn}}^{\tilde{i}}$$(\boldsymbol{x},\boldsymbol{y})\sim D_{\textrm{trn}}^{i}$eward {}$\mathcal{R}_{\{\cdot\}}(i)$
Figure 1: The overview of Mixture-of-Skills. The training collection $D_{\textrm{trn}}=\{D_{\textrm{trn}}^{i}\}_{i=1}^{N}$ consists of various SFT datasets, with $D_{\textrm{trn}}^{i}$ indicating the $i$-th dataset. Please refer to Section 3 for more details.
One common challenge in fine-tuning models across multiple datasets is dealing with their heterogeneity (where different datasets exhibit distinct characteristics) and imbalance (where datasets vary in size and accessibility), making the fine-tuning process highly challenging. To address this challenge, recent approaches often cap dataset usage to prevent models from being overwhelmed by large datasets, but this limits the utilization of all available data (Raffel et al., 2020; Chung et al., 2022; Iyer et al., 2022; Wei et al., 2022). Previous multilingual research adjusts dataset distribution heuristically with a temperature term $\tau$, which requires extensive hyperparameter tuning and overlooks dataset interactions and model learning dynamics (Arivazhagan et al., 2019; Conneau et al., 2020). This leads us to a critical research question: Is there a better way to optimize the data usage?
Building on the research question posed, we first confirm that adjusting the dataset usage properly can significantly enhance model performance (see Table 2). Moreover, inspired by Differentiable Data Selection (Wang et al., 2020a), we propose a general, model-agnostic reinforcement learning framework Mixture-of-Skills (MoS) that learns to optimize data usage automatically during the fine-tuning process. To achieve this optimization, we introduce another set of parameters, $\boldsymbol{\psi}$, known as the scorer network. As shown in Figure 1, this network dynamically adjusts data usage based on the current learning state of the LLM $\boldsymbol{\theta}$. Furthermore, the rewards used to update the scorer network $\boldsymbol{\psi}$ are provided by the LLM $\boldsymbol{\theta}$ from three distinct perspectives: transferability, difficulty, and learning trajectory, ensuring that the scorer network can effectively guide the data usage optimization process. All these efforts constitute the success of our Mixture-of-Skills framework.
To validate the effectiveness of MoS, we conduct extensive experiments using three diverse model backbones: Qwen1.5-0.5B (Bai et al., 2023), Gemma-2B (Mesnard et al., 2024), and Llama-3-8B (AI@Meta, 2024), on two widely-used benchmarks: MMLU (Hendrycks et al., 2021a) and MT-bench (Zheng et al., 2023). Our empirical results demonstrate that MoS substantially improves the models’ overall performance. Our analysis indicates that our model not only effectively learns optimal data utilization but also accelerates training convergence by $\times2$. Additionally, it demonstrates robustness against variations in sampling priors and integrates seamlessly with advanced instance selection methods. Furthermore, we explore the application of MoS in task-specific fine-tuning. We show that MoS, with minor reward modifications known as MoSpec, can be effectively used for task-specific fine-tuning.
Our contributions are summarized as follows:
  • We present a general, model-agnostic reinforcement learning framework, Mixture-of-Skills (MoS), that learns to automatically optimize data usage during the SFT process with three novel rewards (see Section 3).
  • Extensive experiments with three model backbones on two benchmarks demonstrate that MoS significantly enhances model performance. Our analysis reveals that MoS not only effectively learns the optimal data usage but also accelerates training convergence by $\times2$. Additionally, it maintains robustness against variations in sampling priors and is compatible with strong instance selection methods (see Section 4 and Section 5).
  • We explore the application of MoS in task-specific fine-tuning, introducing a variant called MoSpec. This variant, with minor modifications to the rewards, is proven to effectively harness diverse datasets for task-specific fine-tuning (see Section 6).
2 Preliminaries
Supervised Fine-Tuning
A large language model (LLM), parameterized by $\boldsymbol{\theta}$, is capable of following and responding to human instructions after supervised fine-tuning (SFT). Given a single training dataset $D_{\textrm{trn}}^{1}=\{(\boldsymbol{x}_{j},\boldsymbol{y}_{j})\}_{j=1}^{M_{1}}$, where $M_{1}$ is the size of $D_{\textrm{trn}}^{1}$ and $\boldsymbol{x}_{j}$ and $\boldsymbol{y}_{j}$ are the instruction and response of the $j$-th example, the objective function during SFT is to minimize the negative log-likelihood with respect to $\boldsymbol{\theta}$:
$\mathcal{L}_{s}(D_{\textrm{trn}}^{1};\boldsymbol{\theta})=-\sum^{M_{1}}_{j=1}\log p(\boldsymbol{y}_{j}|\boldsymbol{x}_{j};\boldsymbol{\theta}).$ (1)
When fine-tuning the LLM $\boldsymbol{\theta}$ over multiple datasets $D_{\textrm{trn}}=\{D_{\textrm{trn}}^{i}\}_{i=1}^{N}$, where $D_{\textrm{trn}}^{i}=\{(\boldsymbol{x}^{i}_{j},\boldsymbol{y}^{i}_{j})\}_{j=1}^{M_{i}}$, the objective function becomes:
$\mathcal{L}(D_{\textrm{trn}};\boldsymbol{\theta})=\sum^{N}_{i=1}\mathcal{L}_{s}(D_{\textrm{trn}}^{i};\boldsymbol{\theta}).$ (2)
Heuristic Balancing by Temperature
Instead of merging all datasets into a single training mixture, a common approach is to adjust the sampling probability of texts in different languages using a temperature term $\tau$ (Arivazhagan et al., 2019; Conneau et al., 2020). Specifically, the sampling probability of the $i$-th dataset is $q(i)=\frac{|M_{i}|}{\sum^{N}_{n=1}|M_{n}|}$ and can be adjusted by the temperature $\tau$ as:
$q_{\tau}(i)=\frac{q(i)^{1/\tau}}{\sum_{n=1}^{N}q(n)^{1/\tau}}.$ (3)
Consequently, $\tau=1$ corresponds to proportional sampling, equivalent to Equation 2. Conversely, $\tau=\infty$ corresponds to uniform sampling, where smaller datasets are up-sampled to match the largest dataset. The loss function becomes:
$\mathcal{L}(D_{\textrm{trn}};\boldsymbol{\theta},q_{\tau}(N))=\mathbb{E}_{i\sim q_{\tau}(N)}\left[\mathcal{L}_{s}(D_{\textrm{trn}}^{i};\boldsymbol{\theta})\right].$ (4)
Differentiable Data Selection (DDS)
Wang et al. (2020a) propose a general framework that automatically re-weighs training instances to enhance model performance, utilizing a validation set $D_{\textrm{dev}}$. This framework consists of two components: the model $\boldsymbol{\theta}$ and the scorer network $\boldsymbol{\psi}$. The scorer network $\boldsymbol{\psi}$ is designed to calculate a sampling probability for each training instance, which reflects its impact on validation performance. Training instances that have a greater similarity with $D_{\textrm{dev}}$ are allocated a higher probability, thus increasing their likelihood of being selected for model $\boldsymbol{\theta}$ updates. MoS is inspired by DDS but has key differences. Firstly, MoS focuses on rebalancing datasets, unlike DDS, which reweighs training instances. Secondly, MoS does not require prior knowledge of downstream applications, whereas DDS relies on validation set feedback, risking overfitting to that specific validation set. Thirdly, DDS uses the same architecture for the scorer network $\boldsymbol{\psi}$ and model $\boldsymbol{\theta}$, limiting its scalability, while MoS uses a simple MLP model as its scorer network (see Section 3.1).
3 Mixture-of-Skills
In this section, we provide a detailed overview of Mixture-of-Skills (MoS). We begin by outlining the reinforcement learning framework employed in MoS (Section 3.1). Following this, we discuss the reward functions (Section 3.2).
Input : $D_{\textrm{trn}}=\{\{(\boldsymbol{x}_{j}^{i},\boldsymbol{y}_{j}^{i})\}^{M_{i}}_{j=1}\}^{N}_{i=1}$, $N$ training datasets with the size of $M_{i}$ for the $i$-th dataset; $S$, update frequency of $\boldsymbol{\psi}$; $T$, total training steps; $\alpha$, learning rate for $\boldsymbol{\theta}$; $\gamma$, learning rate for $\boldsymbol{\psi}$;
Output : The converged model $\boldsymbol{\theta}$;
1 Initialize $p_{\boldsymbol{\psi}_{0}}(N)$ as Equation 3 with $\tau=\infty$;
2 for t=0 to T do
3       $\tilde{i}\sim p_{\boldsymbol{\psi}}(N)$;
4       Sample batch $(\boldsymbol{x},\boldsymbol{y})\sim D_{\textrm{trn}}^{\tilde{i}}$;
5       $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\alpha\cdot\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{y}|\boldsymbol{x};\boldsymbol{\theta})$;
6       if t % $S$ == 0 then
7             for i=1 to N do
8                   $(\boldsymbol{x}^{\prime},\boldsymbol{y}^{\prime})\sim D_{\textrm{trn}}^{i}$;
9                   Compute reward {}$\mathcal{R}_{\{\cdot\}}(i)$ for $D_{\textrm{trn}}^{i}$ as in Section 3.2;
10
11             end for
12            $\boldsymbol{\psi}\leftarrow\boldsymbol{\psi}+\sum_{i=1}^{N}\gamma\cdot\mathcal{R}_{\{\cdot\}}(i)\cdot\nabla_{\boldsymbol{\psi}}\log p_{\boldsymbol{\psi}}(i)$
13       end if
14
15 end for
Algorithm 1 Mixture-of-Skills
3.1 Learning to Optimize Data Usage
We propose Mixture-of-Skills (MoS) that learns to optimize the data usage during the fine-tuning process by training a scorer network, parameterized by $\boldsymbol{\psi}$, within a reinforcement learning (RL) framework. In this setup, the LLM $\boldsymbol{\theta}$ and the training datasets $D_{\textrm{trn}}$ constitute the environment, while our scorer network serves as the RL agent. In this framework, unlike the static sampling probabilities described in Equation 3, the scorer network $\boldsymbol{\psi}$ dynamically adjusts the sampling probabilities for each dataset in $D_{\textrm{trn}}$ according to the current learning state of the LLM $\boldsymbol{\theta}$. Alternately, the LLM $\boldsymbol{\theta}$ is optimized based on the sampling distribution given by the scorer network $\boldsymbol{\psi}$.
To provide a broader perspective, MoS can be conceptualized as the resolution of a bi-level optimization problem (Colson et al., 2007). In this view, the outer level optimizes the parameters of the LLM $\boldsymbol{\theta}$, while the inner level focuses on adjusting the sampling probabilities through the scorer network $\boldsymbol{\psi}$. Hence, the training objective becomes:
$\boldsymbol{\psi}$ $=\operatorname*{argmin}_{\boldsymbol{\psi}}{\mathcal{J}(D_{\textrm{trn}};\boldsymbol{\theta}(\boldsymbol{\psi}))},\text{where}$ (5)
$\boldsymbol{\theta}(\boldsymbol{\psi})$ $=\operatorname*{argmin}_{\boldsymbol{\theta}}\mathbb{E}_{i\sim p_{\boldsymbol{\psi}}(N)}[\mathcal{L}(D_{\textrm{trn}}^{i};\boldsymbol{\theta})].$
Specifically, we present the algorithm of MoS in Algorithm 1. MoS initially parameterizes the initial sampling probability distribution with $\boldsymbol{\psi}$ as shown in Equation 3, using a warm-up temperature $\tau=\infty$ (see Algorithm 1). When updating the LLM $\boldsymbol{\theta}$, we employ the standard gradient-based optimization method (see Algorithm 1). For computational efficiency, the scorer network $\boldsymbol{\psi}$ is updated every $S$ steps (see Algorithm 1). During updates of $\boldsymbol{\psi}$, we randomly draw one mini-batch from each training set $\{D_{\textrm{trn}}^{i}\}_{i=1}^{N}$ and compute the corresponding rewards as described in Section 3.2 (see Algorithm 1). The training dataset $D_{\textrm{trn}}^{i}$ that yields a high reward is considered to be relatively more beneficial to the overall performance, and its corresponding sampling probability is increased (see Algorithm 1).
A critical issue in Algorithm 1 is that Equation 5 is not directly differentiable with respect to $\boldsymbol{\psi}$. To address this, reinforcement learning (RL) with suitable reward functions is needed Wang et al. (2020a). The update rule for $\boldsymbol{\psi}$ becomes:
$\boldsymbol{\psi}\leftarrow\boldsymbol{\psi}+\sum_{i=1}^{N}\mathcal{R}_{\{\cdot\}}(i)\cdot\nabla_{\boldsymbol{\psi}}\log p_{\boldsymbol{\psi}}(i).$ (6)
Details for the rewards {}$\mathcal{R}_{\{\cdot\}}(i)$ are in Section 3.2 and the update of the scorer network $\boldsymbol{\psi}$ follows the REINFORCE algorithm (Williams, 1992).
Implementing the scorer network
Given that the scorer network $\boldsymbol{\psi}$ is primarily designed to model a relatively simple distribution over the training datasets $D_{\textrm{trn}}$, we utilize a fully connected 2-layer perceptron network for this task. The network takes as input a vector that specifies which training datasets are accessible. Note that the scorer network $\boldsymbol{\psi}$ is used for adjusting the data usage during the fine-tuning process and is orthogonal to the reward model used in RLHF.
3.2 Rewards for Learning
We design the rewards of MoS from three perspectives: transferability (Section 3.2.1), difficulty (Section 3.2.2), and learning trajectory (Section 3.2.3).
3.2.1 Transferability
Transferring knowledge from one problem to another related problem is beneficial, and this transferability is often measured by the similarity between datasets (Du et al., 2018; Zhuang et al., 2021). Datasets with higher similarity are more likely to make more contributions to the targeted performance of the model.
In this work, we represent the training datasets $D_{\textrm{trn}}$ using the mini-batch embeddings and then calculate the pairwise cosine similarities among the mini-batch embeddings for each dataset. We draw a random mini-batch $B^{i}=\{(\boldsymbol{x}^{i}_{j},\boldsymbol{y}^{i}_{j})\}_{j=1}^{L}$ from $D_{\textrm{trn}}^{i}$, where $L$ is the batch size, and the mini-batch embedding $\boldsymbol{z}^{i}$ is defined as:
$\boldsymbol{z}^{i}=\frac{1}{L}\sum^{L}_{j=1}\boldsymbol{e}^{i}_{j},\quad\boldsymbol{e}^{i}_{j}=\frac{1}{K}\sum^{K}_{k=1}\boldsymbol{h}_{k},$ (7)
where $K$ is the sequence length of the concatenation of $\boldsymbol{x}^{i}_{j}$ and $\boldsymbol{y}^{i}_{j}$, and $\boldsymbol{h}_{k}$ is the hidden state of the token $w_{k}$ in the concatenated sequence from the topmost layer of the LLM $\boldsymbol{\theta}_{t}$. Consequently, we define the reward $\mathcal{R}_{\textsc{CosSim}}(i)$ for $D_{\textrm{trn}}^{i}$ as the average cosine similarity among all training datasets:
$\mathcal{R}_{\textsc{CosSim}}(i)=\frac{1}{N}\sum^{N}_{n=1}\frac{\boldsymbol{z}^{i}\cdot\boldsymbol{z}^{n}}{\left\lVert\boldsymbol{z}^{i}\right\rVert\cdot\left\lVert\boldsymbol{z}^{n}\right\rVert},$ (8)
where $N$ is the number of datasets in $D_{\textrm{trn}}$.
3.2.2 Difficulty
Recent work demonstrates that the transfer of knowledge between datasets is not always guaranteed (Wu et al., 2021). In response, we attempt to design the reward based on the inherent difficulty of the dataset in this section.
Recently, the perplexity is used for measure the dataset difficulty (Li et al., 2023b; Marion et al., 2023). Given a training example $(\boldsymbol{x}^{i}_{j},\boldsymbol{y}^{i}_{j})$ from $D_{\textrm{trn}}^{i}$ and the LLM $\boldsymbol{\theta}$, the perplexity is defined as:
PPL $(\boldsymbol{y}^{i}_{j};\boldsymbol{x}^{i}_{j},\boldsymbol{\theta})$ (9)
$=$ $\text{exp}\biggl{(}-\frac{1}{|\boldsymbol{y}^{i}_{j}|}\sum_{k=1}^{|\boldsymbol{y}^{i}_{j}|}\text{log}p_{\boldsymbol{\theta}}(y_{j,k}|\boldsymbol{x}^{i}_{j},\boldsymbol{y}_{j,<k})\biggl{)}.$
However, we argue that perplexity is not a suitable metric for evaluating the difficulty of non-natural language texts, such as mathematical formulas and programming codes. Our preliminary study indicates that the perplexity scores given to mathematical texts by various language models are typically lower than those for natural language texts, despite the common belief that mathematical problems pose significant challenges for LLMs (Yue et al., 2023; Yu et al., 2023). Our preliminary study is presented in Appendix A. Therefore, given a random mini-batch $B^{i}=\{(\boldsymbol{x}^{i}_{j},\boldsymbol{y}^{i}_{j})\}_{j=1}^{L}$ from $D_{\textrm{trn}}^{i}$, the reward $\mathcal{R}_{\textsc{Diff}}(i)$ for $D_{\textrm{trn}}^{i}$ is:
$\mathcal{R}_{\textsc{Diff}}(i)=\frac{1}{L}\sum^{L}_{j=1}\frac{\text{PPL}(\boldsymbol{y}^{i}_{j};\boldsymbol{x}^{i}_{j},\boldsymbol{\theta})}{\text{PPL}(\boldsymbol{y}^{i}_{j};\boldsymbol{x}^{i}_{j},\boldsymbol{\theta}_{0})},$ (10)
where $\boldsymbol{\theta}_{0}$ is the original LLM backbone and $\boldsymbol{\theta}$ is the fine-tuned LLM. The term $\mathcal{R}_{\textsc{Diff}}(i)$ represents the relative decrease in perplexity for $D_{\textrm{trn}}^{i}$ after fine-tuning. A high value of $\mathcal{R}_{\textsc{Diff}}(i)$ suggests that $D_{\textrm{trn}}^{i}$ is difficult to learn and requires more training efforts, while a lower value indicates the opposite.
3.2.3 Learning Trajectory
We design the rewards $\mathcal{R}_{\textsc{CosSim}}(i)$ and $\mathcal{R}_{\textsc{Diff}}(i)$ based on the transferability and difficulty of the training dataset $D_{\textrm{trn}}^{i}$, as discussed in Section 3.2.1 and Section 3.2.2. However, both rewards ignore the learning trajectory of the fine-tuning process. Therefore, we introduce the exponential moving average (EMA) when estimating the rewards. This approach can both better estimate the reward and stabilize the data usage optimization process. Specifically, we define the EMA as follows:
{}$\mathcal{R}_{\{\cdot\}}(i)=\beta\mathcal{R}^{{}^{\prime}}_{\{\cdot\}}(i)+(1-\beta)\mathcal{R}^{{}^{\prime\prime}}_{\{\cdot\}}(i),$ (11)
where $\beta$ is the smoothing factor, {}$\mathcal{R}^{{}^{\prime}}_{\{\cdot\}}(i)$ indicates the original reward for the current update, {}′′$\mathcal{R}^{{}^{\prime\prime}}_{\{\cdot\}}(i)$ represents the reward for the previous update, and {}$\mathcal{R}_{\{\cdot\}}(i)$ is the smoothed reward for the current update. Note that both $\mathcal{R}_{\textsc{CosSim}}(i)$ and $\mathcal{R}_{\textsc{Diff}}(i)$ can be applied in Equation 11 and we set $\beta=0.9$.
4 Experiments
We present our experimental setup (Section 4.1) and main results (Section 4.2) in this section.
4.1 Experimental Setup
#exam. #words Inst.L Resp.L Turns
Mathematics 026.2K 03.4M 047.6 084.0 1.0
Medicine 005.2K 01.0M 036.5 147.4 1.0
General 009.3K 09.3M 054.3 202.4 3.6
NLP 062.6K 08.6M 127.9 009.7 1.0
Total 103.4K 22.3M 088.4 084.3 1.2
Table 1: Dataset statistics of the training datasets in this work. Inst.L, Resp.L, and Turns indicate the average of instruction length (in words), response length (in words), and number of conversation turns.
Datasets
In this work, we collect four distinct supervised fine-tuning (SFT) datasets:
Due to the constraint on compute, we sample 10% of examples from each dataset and present the dataset statistics in Table 1.
$\mu_{\textrm{BOTH}}$ MMLU MT-bench
$\mu_{\textrm{MU}}$ Math Med. Others $\mu_{\textrm{MB}}$ Turn 1 Turn 2
Qwen1.5-0.5B
Prop. ($\tau=1$) 32.82 30.95 23.40 31.30 31.76 3.47 4.18 2.76
Temp. ($\tau=10$) 34.17 32.09 22.88 30.88 33.41 3.63 4.11 3.14
Uni. ($\tau=\infty$) 33.81 31.52 21.45 29.97 33.02 3.61 4.21 3.01
[…]
MultiUAT 34.15 31.24 21.41 29.91 33.01 3.71 4.26 3.15
[…]
MoS + CosSim + EMA 35.13 32.45 22.27 31.56 33.82 3.78 4.44 3.13
MoS + Diff 34.24 31.49 20.71 29.94 33.07 3.70 4.21 3.19
MoS + Diff + EMA 34.83 32.11 21.84 31.01 33.53 3.76 4.11 3.40
Gemma-2B
Prop. ($\tau=1$) 42.90 33.61 20.03 33.33 35.25 5.22 5.63 4.81
Temp. ($\tau=10$) 41.86 36.16 21.03 37.92 37.55 4.76 5.49 4.03
Uni. ($\tau=\infty$) 43.95 35.95 20.82 35.97 37.71 5.19 5.60 4.79
[…]
MultiUAT 43.19 33.95 20.85 34.64 33.71 5.24 5.61 4.88
[…]
MoS + CosSim + EMA 44.49 33.86 20.57 33.32 35.52 5.51 6.01 5.01
MoS + Diff 44.93 34.32 20.29 34.06 36.00 5.55 6.04 5.06
MoS + Diff + EMA 45.10 34.61 20.83 34.64 36.20 5.56 6.13 4.99
Llama-3-8B
Prop. ($\tau=1$) 60.97 56.78 26.61 62.03 59.19 6.52 6.96 6.08
Temp. ($\tau=10$) 61.40 56.17 28.36 59.64 58.68 6.66 7.04 6.29
Uni. ($\tau=\infty$) 60.99 55.72 27.65 60.77 57.93 6.63 7.04 6.21
[…]
MultiUAT 61.18 55.66 28.65 60.48 57.32 6.67 7.05 6.29
[…]
MoS + CosSim + EMA 63.85 58.08 27.60 61.54 60.90 6.96 7.28 6.65
MoS + Diff 63.00 57.93 31.08 62.65 60.07 6.81 6.98 6.64
MoS + Diff + EMA 63.26 58.34 32.81 62.21 60.49 6.82 7.34 6.30
Table 2: Main results given by Qwen1.5-0.5B, Gemma-2B, and Llama-3-8B on MMLU and MT-bench. The best and second-best results are highlighted in bold and underline. Note that $\mu_{\textrm{MB}}$ is upscaled by 10×10$\times$10 × to a range from 1 to 100 used for computing $\mu_{\textrm{BOTH}}$.
Model Backbones
We apply MoS to three diverse model backbones, including Qwen1.5-0.5B (Bai et al., 2023), Gemma-2B (Mesnard et al., 2024), and Llama-3-8B (AI@Meta, 2024).
Optimization
We fine-tune all the parameters of large language models (LLMs) using the AdamW optimizer (Kingma and Ba, 2015; Loshchilov and Hutter, 2019) with a learning rate of $1\times 10^{-5}$ and a batch size of 64. We fine-tune all the models for 3 epochs, or the equivalent number of steps. During the training process, we apply the linear learning rate schedule, which includes a warm-up phase comprising 10% of the total training steps. For MoS, $\boldsymbol{\psi}$ is updated for every 100 steps with the learning rate of $1\times 10^{-4}$ and the batch size of 64. $\boldsymbol{\psi}$ is initialized by $\tau=\infty$ in Equation 3.
Baselines
We compare MoS with several heuristic and dynamic baselines as follows:
  • Heuristic: Based on Equation 3, we consider proportional sampling (Prop., $\tau=1$), temperature sampling (Temp., $\tau=10$), and uniform sampling (Uni., $\tau=\infty$).
  • Dynamic: MultiDDS (Wang et al., 2020b) and MultiUAT (Wu et al., 2021) dynamically balance the dataset distribution using the gradient cosine similarity and model uncertainty on the validation sets as rewards, respectively. We sample 1K examples from each dataset as the validation sets for both approaches.
We do not include maximum cap and other instance selection methods as baselines because they do not fully utilize all available data.
Evaluation
In this work, we evaluate two widely-used benchmarks that are highly correlated with human judgments:
  • MMLU: Hendrycks et al. (2021a) propose the MMLU benchmark, covering 57 subjects across STEM, humanities, social sciences, and more. We categorize the subjects into three groups: mathematics, medicine, and others, and conduct zero-shot evaluations. We report the average accuracy for each group and the accuracy across all subjects, denoted as $\mu_{\textrm{MU}}$. Detailed subject categorization is in Appendix B.
  • MT-bench: Zheng et al. (2023) propose the MT-bench, a multi-turn conversational benchmark designed to measure large language models’ capabilities. This benchmark covers eight key skills, including coding, writing, roleplay, and more. LLM responses are scored by gpt-4 on a scale from 1 to 10. The overall score across all eight skills is denoted as $\mu_{\textrm{MB}}$.
The overall performance is reported as the average score of both $\mu_{\textrm{MU}}$ and $\mu_{\textrm{MB}}$, denoted as $\mu_{\textrm{BOTH}}$. Note that when computing $\mu_{\textrm{BOTH}}$, MT-bench scores are upscaled by 10×10$\times$10 × to range from 1 to 100, maintaining consistency with MMLU.
4.2 Main Results
We present the main results in Table 2.
An optimal temperature $\tau$ boosts performance, but no universally optimal $\tau$ exists.
When comparing the heuristic baselines, we observe that there is no universally optimal $\tau$ that consistently works well for all model backbones. As shown in Table 2, Temp. ($\tau=10$) performs best for Qwen1.5-0.5B and Llama-3-8B, but is least effective for Gemma-2B. This variability confirms the motivation behind this work.
MoS outperforms heuristic baselines, with larger models showing greater improvements.
Our method consistently outperforms heuristic baselines across all three model backbones in terms of $\mu_{\textrm{BOTH}}$. Notably, larger models show greater improvements with our approach. As shown in Table 2, the best variant of our method surpasses the best heuristic baseline by +0.96, +1.15, and +2.45 in $\mu_{\textrm{BOTH}}$ for Qwen1.5-0.5B, Gemma-2B, and Llama-3-8B, respectively. This is particularly significant in the era of LARGE language models.
Different rewards work better for different models, and EMA always helps.
As shown in Table 2, MoS with CosSim outperforms MoS with Diff for Qwen1.5-0.5B and Llama-3-8B, while Diff-based MoS yields better results for Gemma-2B. Additionally, EMA consistently enhances overall performance in terms of $\mu_{\textrm{BOTH}}$, supporting our rationales concerning learning trajectory in Section 3.2.3.
005005005005001,00010001{,}0001 , 0001,50015001{,}5001 , 5002,00020002{,}0002 , 0002,50025002{,}5002 , 5003,00030003{,}0003 , 0003,50035003{,}5003 , 5004,00040004{,}0004 , 0004,50045004{,}5004 , 5005,00050005{,}0005 , $0$.200.200.200.200.250.250.250.250.300.300.300.300.350.350.350.35Trainning StepsSamp. Prob.MathematicsMedicineGeneralNLP
(a) MoS + CosSim + EMA
005005005005001,00010001{,}0001 , 0001,50015001{,}5001 , 5002,00020002{,}0002 , 0002,50025002{,}5002 , 5003,00030003{,}0003 , 0003,50035003{,}5003 , 5004,00040004{,}0004 , 0004,50045004{,}5004 , 5005,00050005{,}0005 , $0$.200.200.200.200.250.250.250.250.300.300.300.300.350.350.350.35Trainning StepsSamp. Prob.
(b) MoS + Diff + EMA
Figure 2: Learned dataset distribution given by Llama-3-8B with different variations of MoS. The $x$-axis indicates the training steps, and the $y$-axis indicates the sampling probabilities of datasets.
5 Analysis
In this section, we conduct an in-depth analysis of MoS using Llama-3-8B. Our analysis encompasses the learned dataset distribution of MoS, the training convergence speed of MoS, the impact of sampling priors on MoS, and its compatibility with the instance selection methods.
MoS with different rewards learns different dataset distributions.
We visualize the dataset distribution learned by MoS using Llama-3-8B as shown in Figure 2. Starting with equal sampling probabilities, both CosSim and Diff variations of MoS increase the probability for General and decrease it for Medicine. However, CosSim maintains the probabilities for Mathematics and NLP, whereas Diff upsamples Mathematics but downsamples NLP.
001,00010001{,}0001 , $0002{,}$ $0003{,}$ $0004{,}$ $0005{,}$ $0$.200.200.200.200.400.400.400.400.600.600.600.600.800.800.800.802.2×$\times2$ fasterTraining StepsTraining LossProp. ($\tau=1$)Temp. ($\tau=10$)Uni. ($\tau=\infty$)MoS + Diff + EMA
Figure 3: Training loss curves of heuristic baselines and MoS + Diff + EMA.
MoS speeds up the convergence.
We illustrate the training dynamics of heuristic baselines and MoS + Diff + EMA in Figure 3. When compared to Temp. ($\tau=10$), which is the best-performing heuristic baseline, MoS + Diff + EMA demonstrates notable improvements. Specifically, it converges approximately $\times2$ faster, as shown in Figure 3, and achieves a +1.86 improvement in terms of $\mu_{\textrm{BOTH}}$, as detailed in Table 2.
$\mu_{\textrm{BOTH}}$ $\mu_{\textrm{MU}}$ $\mu_{\textrm{MB}}$
Temp. ($\tau=10$) 61.40 56.17 6.66
MoS + CosSim + EMA
    + $\tau=1$ 62.49 56.95 6.80
    + $\tau=10$ 63.94 57.98 6.99
    + $\tau=\infty$ 63.85 58.08 6.96
MoS + Diff + EMA
    + $\tau=1$ 62.29 56.51 6.81
    + $\tau=10$ 63.66 58.22 6.91
    + $\tau=\infty$ 63.26 58.34 6.82
Table 3: Results of MoS with different sampling priors $\tau$. The best results are highlighted in bold.
MoS demonstrates robustness to changes in sampling priors.
As indicated in Algorithm 1 in Algorithm 1, we initialize our sampling probability distribution with $\tau=\infty$. Consequently, we investigate the effects of various sampling priors on MoS. As shown in Table 3, MoS with different sampling priors consistently outperforms the best heuristic baseline Temp. ($\tau=10$). Moreover, selecting the appropriate sampling prior for MoS can further enhance performance. These results underscore the robustness and effectiveness of our approach.
$\mu_{\textrm{BOTH}}$ $\mu_{\textrm{MU}}$ $\mu_{\textrm{MB}}$
Random
Prop. ($\tau=1$) 59.78 54.39 6.52
Temp. ($\tau=10$) 60.62 54.89 6.63
Uni. ($\tau=\infty$) 59.21 54.21 6.42
IFD
Prop. ($\tau=1$) 60.43 55.03 6.58
Temp. ($\tau=10$) 61.02 55.13 6.69
Uni. ($\tau=\infty$) 60.62 54.92 6.63
[…]
MoS + Diff + EMA 62.05 56.43 6.77
Table 4: Compatibility between MoS and IFD. Random and IFD indicate the 10% data from each dataset selected by random sampling and IFD selection, respectively. The best results are highlighted in bold.
MoS is compatible with the instance selection method.
Following Li et al. (2023b), we leverage Qwen1.5-0.5B to calculate the Instruction-Following Difficulty (IFD) scores for each training instance and select the top 10% of training instances with the highest scores from each dataset. As shown in Table 4, combining MoS with IFD further improves the model performance, indicating the successful combination of MoS and IFD.
6 Fine-tuning from Generalist to Specialist
$\mu_{\textrm{ALL}}$ GSM8K MATH M-math
5-shot 5-shot 0-shot
Generalist
Temp. ($\tau=10$) 29.17 49.62 09.54 28.36
MoS + CosSim + EMA 29.26 50.40 09.78 27.60
MoS + Diff + EMA 30.88 49.58 10.26 32.81
Math Specialist
MathLlama-3-8B 27.04 41.02 09.76 30.34
MoSpec + C.S. + E. 30.63 51.10 10.64 30.16
MoSpec + D. + E. 31.91 52.10 11.40 32.24
Table 5: Results on GSM8K, MATH, and M-math given by generalists and math specialists. $\mu_{\textrm{ALL}}$ indicates the average performance over all three benchmarks. MoSpec + C.S. + E. and MoSpec + D. + E. indicates MoSpec + CosSim + EMA and MoSpec + Diff + EMA, respectively. The best results are highlighted in bold.
Large, general-purpose models offer broad capabilities but can be costly to deploy in real-world applications. Many scenarios require only a narrow set of functionalities, making smaller, specialized models more effective for specific tasks than larger, general-purpose ones (Luo et al., 2023; Azerbayev et al., 2023; Wu et al., 2024a). Mixture-of-Skills (MoS) is a framework designed to optimize data usage for various fine-tuning purposes, including task-specific fine-tuning. This section explores the application of MoS in this context.
We aim to fine-tune Llama-3-8B to specialize in mathematics using datasets from Section 4.1, referring to this modified version as MoSpec. For MoSpec with CosSim, we compute the cosine similarity between Mathematics and other datasets, including Mathematics itself. For MoSpec with Diff, we double the reward for the Mathematics dataset. For comparison, we fine-tune Llama-3-8B directly on Mathematics dataset in Section 4.1, denoted as MathLlama-3-8B. Both MoSpec and MathLlama-3-8B use the identical hyperparameters from Section 4.1, except MathLlama-3-8B is fine-tuned for 12 epochs for fairness. We evaluate the models on math-related subjects in MMLU (0-shot, denoted as M-math), GSM8K (5-shot) (Cobbe et al., 2021), and MATH (5-shot) (Hendrycks et al., 2021b).
SFT datasets from other sources are beneficial for the specific target task.
As shown in Table 5, the MathLlama-3-8B model trained solely on the Mathematics subset performs the worst among all models. This indicates that incorporating additional SFT datasets is advantageous. The performance gap is particularly evident on the GSM8K dataset, which requires step-by-step reasoning. We believe this discrepancy arises from the Mathematics subset’s incompleteness, while other SFT datasets can address these shortcomings.
005005005005001,00010001{,}0001 , 0001,50015001{,}5001 , 5002,00020002{,}0002 , 0002,50025002{,}5002 , 5003,00030003{,}0003 , 0003,50035003{,}5003 , 5004,00040004{,}0004 , 0004,50045004{,}5004 , 5005,00050005{,}0005 , $0$.100.100.100.100.200.200.200.200.300.300.300.300.400.400.400.40Trainning StepsSamp. Prob.MathematicsMedicineGeneralNLP
005005005005001,00010001{,}0001 , 0001,50015001{,}5001 , 5002,00020002{,}0002 , 0002,50025002{,}5002 , 5003,00030003{,}0003 , 0003,50035003{,}5003 , 5004,00040004{,}0004 , 0004,50045004{,}5004 , 5005,00050005{,}0005 , $0$.100.100.100.100.200.200.200.200.300.300.300.300.400.400.400.40Trainning StepsSamp. Prob.
Figure 4: Learned dataset distribution given by Llama-3-8B with MoSpec + CosSim + EMA (left) and MoSpec + Diff + EMA (right).
MoSpec can harness diverse datasets to enhance task-specific performance.
When comparing MoS and MoSpec with the same reward type, MoSpec consistently outperforms MoS on mathematical benchmarks in Table 5. By assigning a higher reward, MoSpec effectively learns the optimal dataset distribution for learning the mathematical capabilities. As shown in Figure 4, CosSim and Diff in MoSpec effectively guide the scorer network $\boldsymbol{\psi}$ to increase the sampling probability of Mathematics.
We believe this property of MoS is particularly meaningful when the task-specific dataset is not sufficiently large or comprehensive.
7 Related Work
Data Engineering for LLMs
The success of large language models (LLMs) heavily relies on their training datasets. Researchers gather or create extensive datasets (Raffel et al., 2020; Gao et al., 2021; Penedo et al., 2023; Wang et al., 2023; Li et al., 2023a; Cui et al., 2023; Wu et al., 2024b; Wang et al., 2024a). Recent efforts focus on selecting data subsets to enhance training efficiency. Xie et al. (2023) estimate the quality of each subset in the pretraining dataset mixture using a small proxy model. Recent approaches filter out low-quality examples using perplexity (Li et al., 2023b; Marion et al., 2023).
Dataset Rebalancing
The standard practice for dataset rebalancing in fine-tuning large language models (LLMs) involves capping the number of examples per dataset (Raffel et al., 2020; Chung et al., 2022; Iyer et al., 2022; Wei et al., 2022). However, this approach does not fully utilize all available data. Previous multilingual research often rebalances datasets for multiple languages using a temperature term $\tau$ (Arivazhagan et al., 2019; Conneau et al., 2020). Furthermore, Wang et al. (2020a) reweigh training examples based on their similarity with the validation set. Inspired by Wang et al. (2020a), Wang et al. (2020b) and Wu et al. (2021) propose rebalancing the dataset distribution for machine translation tasks.
Multi-Task Learning
Our work is also related to multi-task learning (Ruder, 2017; Crawshaw, 2020; Zhang et al., 2023b). Both transferability and difficulty are commonly used for reweighting the importance of tasks to achieve better overall performance and mitigate the conflicts between tasks (Kendall et al., 2018; Chen et al., 2018; Yu et al., 2020; Wang et al., 2021). We highlight that tasks are the specific goals the model works towards, while skills are the broader abilities that allow the model to perform a wide range of tasks.
Ours
In this work, Mixture-of-Skills (MoS) is inspired by Wang et al. (2020a) and related to Wang et al. (2020b) and Wu et al. (2021), but offers several key advancements. Unlike previous methods, MoS does not require knowledge of downstream applications, avoiding the risk of overfitting to validation sets. Additionally, MoS introduces novel rewards tailored for LLMs and considers the learning trajectory during fine-tuning, enhancing overall performance. Finally, MoS is highly adaptable for specific fine-tuning needs, setting it apart from prior works.
8 Conclusion
In this work, we address the critical challenge of optimizing data usage during the fine-tuning process of LLMs. We propose a general, model-agnostic reinforcement learning framework, Mixture-of-Skills (MoS), that dynamically adjusts dataset usage to enhance model performance with three novel rewards. Through extensive experiments on three diverse model backbones and two widely-used benchmarks, we demonstrate that MoS significantly improves overall model performance. Additionally, we explore the application of MoS in task-specific fine-tuning, leading to the development of MoSpec. Our experiments show that models fine-tuned with MoSpec on various datasets outperform those trained solely on task-specific datasets. In summary, MoS provides a powerful and flexible solution to the challenges of dataset heterogeneity and imbalance in the fine-tuning of LLMs.
9 Limitations
Computational Overhead
In this study, the scorer network $\boldsymbol{\psi}$ and the large language model (LLM) $\boldsymbol{\theta}$ are updated in an alternating fashion. Although the scorer network $\boldsymbol{\psi}$ is a relatively simple two-layer MLP model, the overall training duration increases by approximately 20%, compared with the heuristic baselines, when the LLM $\boldsymbol{\theta}$ is updated for the same number of steps.
Number of Datasets
Our experiments are limited to four datasets due to computational resource constraints. The performance of our approach as the dataset count increases remains unexplored.
These limitations are acknowledged and we leave them to the future work.
Acknowledgment
This work is partly supported by the ARC Future Fellowship FT190100039. This material is based on research partially supported by the DARPA Assured Neuro Symbolic Learning and Reasoning (ANSR) program under award number FA8750-23-2-1016.
Appendix A Preliminary Study on Perplexity
Perplexity measures how well a probability model predicts a sample, quantifying the model’s uncertainty in making these predictions. It is designed for natural language texts because it relies on the probability distributions typical to human languages. However, non-natural language texts, such as mathematical formulas or programming code, often involve symbols and structures whose relationships are governed by logical or mathematical rules rather than linguistic context. We hypothesize that using perplexity to measure the difficulty in these contexts does not capture the essential aspects of understanding or generating such texts.
To verify our hypothesis, we conduct a preliminary study on perplexity and present the results in Table 6. We observe that the perplexity of Mathematics is commonly lower than that of other datasets given by Qwen1.5-0.5B, Gemma-2B, and Llama-3-8B, regardless of whether the models are fine-tuned, while the perplexity of NLP is the highest among all the datasets. However, Mathematics is associated with a higher value of $\Delta$, while NLP achieves the lowest value of $\Delta$, suggesting that Mathematics is difficult to learn while NLP is easy to learn. If we utilize the perplexity as a measure of difficulty in Section 3.2.2, MoS incorrectly assigns a higher sampling probability to NLP.
Subjects
Mathematics Abstract Algebra, College Mathematics, Elementary Mathematics, High School Mathematics, High School Statistics
Medicine Anatomy, Clinical Knowledge, College Medicine, Human Aging, Human Sexuality, Medical Genetics, Nutrition, Professional Medicine, Virology
Others Astronomy, Business Ethics, College Biology, College Chemistry, College Computer Science, College Physics, Computer Security, Conceptual Physics, Econometrics, Electrical Engineering, Formal Logic, Global Facts, High School Biology, High School Chemistry, High School Computer Science, High School European History, High School Geography, High School Government And Politics, High School Macroeconomics, High School Microeconomics, High School Physics, High School Psychology, High School US History, High School World History, International Law, Jurisprudence, Logical Fallacies, Machine Learning, Management, Marketing, Miscellaneous, Moral Disputes, Moral Scenarios, Philosophy, Prehistory, Professional Accounting, Professional Law, Professional Psychology, Public Relations, Security Studies, Sociology, US Foreign Policy, World Religions
Table 7: MMLU subject categorization.
Appendix B MMLU Subject Categorization
We present the detailed subject categorization of MMLU in Table 7.